OpenStudy (anonymous):

tan α = sin α b. tan α = cos a/sin a c. cos α = 0.5 d. sin α = cos (90° – α)

Parth (parthkohli):

Hint:\[\tan\theta = {\sin\theta \over \cos\theta} \]

OpenStudy (anonymous):

thnks so a would be right

Parth (parthkohli):

no!

OpenStudy (anonymous):

i meant to put b